# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

### v3

@@ 4.1 Experimental Settings
Dataset. We collect literature review articles published in 2023 from the annual reviews website and obtain 1105 articles, the dataset distribution across journals as shown in Appendix B Figure 4.
LLMs Selection. We select five LLMs for evaluation: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct.
In task 1 and task 3, we set $N$ as 10, each model generates 10 references. For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, for Task 2, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; To compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings. For task 3, we employ GPT-4o as the NLI model in Eq 8, and set $q$ as 10.
@@ 4.2 Main Results
Results for Task 1. By analyzing the results in Table 1, we observe that Claude-3.5-Sonnet achieves the highest accuracy and title search rate, whereas Llama-3.2-3B exhibits the lowest performance on both metrics for Task 1. Furthermore, when evaluating the accuracy of LLM-generated references, treating the author item as correct when the first author is identified correctly results in a 3–5% increase in accuracy scores across all models. This finding underscores that accurately generating complete author lists remains a significant challenge for LLMs. Additionally, we conduct a year-wise analysis of the literature that each model cited correctly, as illustrated in the line graph in Figure 2. The graph shows that the majority of the correctly cited references for almost all LLMs involved in the task are concentrated between 2010 and 2020.
Results for Task 2. We use various metrics to evaluate model performance. As shown in Table 2, Claude-3.5-Sonnet consistently achieves the best overall performance across nearly all metrics. The average semantic similarity between the abstracts generated by Claude-3.5-Sonnet and human-written abstracts is 81.17%. Furthermore, Claude-3.5-Sonnet and DeepSeek-V3 exhibit the highest factual consistency, achieving a TURE model evaluation score of 78.10% and a GPT-4o assessment score of 96.84%, respectively. Additionally, we observe that Llama-3.2-3B achieves the highest score on the ROUGE-L metric but does not exhibit a clear advantage on other metrics. This suggests that, given the diversity of outputs produced by LLMs, it is crucial to employ multiple evaluation metrics when assessing their performance on our tasks.
Results for Task 3. As shown in Table 3, we observe that, compared to the accuracy scores in Task 1, all LLMs exhibit a significant increase in accuracy when generating references in Task 3. Previous studies have shown that citing real external sources can reduce hallucination rates in the generated text (Gao et al., 2023). In our experiment, we find that when LLMs are tasked with generating references for the literature review, the accuracy of the generated references increases significantly. This suggests that the generated references and the literature review text impose constraints on each other.
@@
On the other hand, we find that Claude-3.5-Sonnet achieves the best performance on the KPR metric, indicating that the literature review generated by Claude-3.5-Sonnet recalls the most claims from the human-written literature review. While the literature review generated by DeepSeek-V3 performs well on the ROUGE metrics.
  Models Task 1
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ Acc. Acc.(first author)
Qwen-2.5-72B 21.80 12.25 17.58
Llama-3.2-3B 16.62 3.45 6.95
DeepSeek-V3 56.04 46.33 50.66
GPT-4o 32.07 21.65 24.65
Claude-3.5-Sonnet 64.82 51.59 55.77
@@
Table 1: The experimental results of the five LLMs in Task 1. “$S_{t}$” refers to the title search rate as defined in Eq 4, while “Acc.” represents the accuracy defined in Eq 3. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task 2
@@
Table 2: Compare the performance of four LLMs on Task 2.
  Models Task 3
@@
Qwen-2.5-72B 40.02 28.91 33.64 38.82 29.95 9.01 15.14
Llama-3.2-3B 21.78 4.86 7.28 29.07 28.07 7.77 15.46
DeepSeek-V3 62.29 52.81 55.38 56.02 35.65 10.40 17.46
GPT-4o 60.05 50.62 54.16 59.18 30.78 9.72 15.54
Claude-3.5-Sonnet 66.43 59.06 63.06 62.32 28.59 8.90 14.41
Table 3: The experimental results of the four LLMs in Task 3. “$S_{t}$” refers to the title retrieval rate as defined in Eq 4.“Acc.” represents the accuracy defined in Eq 4. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Task1 Task3
Title Acc. Overlap Overlap(first author) Title Acc. Overlap Overlap(first author)
Qwen-2.5-72B 7.20 12.60 13.27 12.15 17.36 18.31
Llama-3.2-3B 5.97 8.48 8.67 8.01 8.28 8.44
DeepSeek-V3 14.28 19.72 20.50 20.44 26.79 27.30
GPT-4o 8.66 18.76 19.50 18.31 27.88 28.86
Claude-3.5-Sonnet 16.06 24.34 25.21 21.20 31.90 33.25
Table 4: Overlap rate of citations between LLM-generated and human-written references in Task 1 and Task 3. “Title Acc.” refers to the title overlap between LLM-generated and human-written references, while “overlap” represents the article overlap between LLM-generated and human-written references. “first author” refers to when evaluating the overlap of references, the author dimension only comparing the first author.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Task 1 and Task 3, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in Task 1. Additionally, the accuracy of reference generation in Task 1 for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Task 3, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4o in certain dimensions approaches or even exceeds that of Claude-3.5-Sonnet. However, the performance of LLaMA-3.2-3B remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
First, we observe that in Task 1, as shown in Figure LABEL:f4a, almost all models exhibit the highest accuracy in the Mathematics discipline and lowest accuracy in the Chemistry discipline. Additionally, we calculate the results of the one-way ANOVA for each LLM across five disciplines to validate the differences. The results are as follows: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065). These results indicate significant differences in performance across disciplines for each model except Llama-3.2-3B.
@@
Secondly, as shown in Figure LABEL:f4b, we present the NLI scores of each model across different disciplines as evaluated by TRUE in Task 2. We observe that all models exhibit the lowest entailment scores in Social science and the highest scores in Biology and Technology. Additionally, we conduct an ANOVA test to evaluate the models’ performance across different disciplines. The results are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001). These results indicate that, in Task 2, there are significant differences in the performance of each model across different disciplines.
@@
Additionally, we examine the references accuracy of each model across five disciplines in Task 3, as illustrated in Figure LABEL:f4c. It is evident that the accuracy of Claude-3.5-Sonnet, DeepSeek-V3, and GPT-4o are significantly higher than those of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines. Furthermore, the trend aligns with that observed in Task 1: almost all models exhibit the highest accuracy in Mathematics, followed by relatively high scores in Social Science, while lower accuracy scores are observed in Chemistry and Technology. The results of the one-way ANOVA for each models across five disciplines are as follows: Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001), indicating significant differences in each models’ performance across the various disciplines.
@@
Finally, we observe the KPR metric across different disciplines in Task 3, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests. Specifically, the p-values are as follows: Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).
4.5 Comparison of LLM-Cited and Human-cited References
In this section, we also compare the LLM-generated references with those from the human-written original articles. Specifically, we examine the overlap between the references generated by each model and those that are cited in the human-written article. As shown in Table 4, we observe that in Task 1, 25.21% of the references generated by Claude-3.5-Sonnet are also cited in the human-written article, indicating a relatively high overlap rate. Furthermore, we find that, compared to Task 1, the reference overlap rate with human citations increases across all models in Task 3. Notably, GPT-4o models show a 10% increase in the overlap rate. We provide a more detailed comparison of the references across various dimensions in Appendix C.
@@ 5 Conclusion
In this paper, we evaluate the ability of LLMs to write literature reviews. However, instead of assessing from conventional perspectives, such as the fluency of the generated context or topic coverage, we compare the model-generated results with human-written ones. Additionally, when processing the LLM-generated results, we often need to handle abbreviations of author names and journal titles. While we have made efforts to address these details thoroughly, some discrepancies may still exist.
@@ Appendix C Comparison of LLM-Cited and Human-cited References from Different Dimensions
We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions. As shown in Figure LABEL:f3, for Task 1, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Task 3, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Task 1. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Task 3 leads to the generation of more accurate author information.
@@ Appendix D Discussion
When comparing the references generated by the models in Task 1 and Task 3, we find that in Task 3, nearly all models generate more accurate references. This suggests that LLMs cite references during the writing process, which improves the authenticity of the references. Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.

### v4

@@ 4.1 Experimental Settings
Dataset. We collect 1,105 literature review articles published in 2023 from the Annual Reviews website. The distribution of articles across journals is shown in Appendix B, Figure 4.
LLMs Selection. We evaluate five LLMs: Claude-3.5-Sonnet-20240620, GPT-4o-2024-08-16, Qwen-2.5-72B-Instruct, DeepSeek-V3, and Llama-3.2-3B-Instruct. All model outputs were generated via their official APIs with temperature set to 0 for consistency.
In Reference Generation and Review Composition tasks, we set $N$ as 10, each model generates 10 references. For the generated reference evaluation, we use Semantic Scholar as the external database. Recently, LLMs-as-judges has become more common Chen et al. (2024); Zheng et al. (2023); Shangyu et al. (2024). So, for the Abstract Writing task, we employ TRUE Honovich et al. (2022), along with GPT-4o as NLI models for evaluating factual consistency in context; to compute semantic similarity, we use text-embedding-3-large to convert texts into embeddings. For the Review Composition task, we employ GPT-4o as the NLI model in Eq 8, and set $q$ as 10.
@@ 4.2 Main Results
Results for Reference Generation. As shown in Table 1, Claude-3.5-Sonnet achieves the highest $F1$ score and $S_{t}$, while Llama-3.2-3B performs the worst on both metrics. Notably, Claude-3.5-Sonnet also achieves the highest precision and recall, indicating that it not only generates more correct references overall but also shares the highest overlap with those cited in the human-written article. When evaluating the author dimension of LLM-generated references, we consider the reference to match in this dimension if the first author is correctly matched. Applying this criterion results in a 1–3% increase in F1 scores across all models. This suggests that generating complete and accurate author lists remains a major challenge for LLMs.
We further conduct a year-wise analysis of the correctly generated references, as illustrated in Figure 2. The results reveal that the majority of accurate citations produced by the models are concentrated in the period between 2010 and 2020, a trend consistent across nearly all LLMs evaluated in this task.
Results for Abstract Writing. As shown in Table 2, Claude-3.5-Sonnet achieves the best overall performance across most evaluation metrics. It generates abstracts with the highest average semantic similarity to human-written ones (81.17%) and shows strong factual consistency, achieving a TRUE score of 78.10%. DeepSeek-V3 also performs well in factual consistency, with the highest GPT-4o-based assessment score (96.84%). In contrast, Llama-3.2-3B obtains the highest ROUGE-L score but does not show clear advantages on other metrics. These results highlight the importance of using multiple evaluation metrics to comprehensively assess the diverse outputs of LLMs.
Results for Review Composition. As shown in Table 3, compared to the Reference Generation task, all LLMs demonstrate a significant increase in Precision when generating references within the Review Composition task. Prior research indicates that grounding generated text with real external citations can effectively reduce hallucination rates (Gao et al., 2023). Consistently, our experiments reveal that when LLMs generate references alongside the literature review, the accuracy of these references improves markedly. This suggests a mutual constraint between the generated references and the review text, leading to enhanced overall reliability.
@@
On the other hand, Claude-3.5-Sonnet achieves the highest performance on the KPR metric, indicating that its generated literature reviews recall the greatest number of claims from the human-written versions. Meanwhile, the literature reviews produced by DeepSeek-V3 excel on the ROUGE metrics, demonstrating stronger overlap with reference texts in terms of lexical similarity.
  Models Reference Generation
$\textbf{{S}}_{\textbf{{t}}}$$\uparrow$ P R F1 P R F1(first author)
Qwen-2.5-72B 21.80 12.25 12.60 12.42 17.58 13.27 15.12
Llama-3.2-3B 16.62 3.45 8.48 4.90 6.95 8.67 7.72
DeepSeek-V3 56.04 46.33 19.72 27.66 50.66 20.50 29.19
GPT-4o 32.07 21.65 18.76 20.10 24.65 19.50 21.77
Claude-3.5-Sonnet 64.82 51.59 24.34 33.08 55.77 25.21 34.72
@@
Table 1: The experimental results of the five LLMs in Reference Generation. “$S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Models Abstract Writing
@@
Table 2: Compare the performance of four LLMs on Abstract Writing.
  Models Review Composition
@@
Qwen-2.5-72B 40.02 28.91 17.36 21.69 33.64 18.31 23.713 38.82 29.95 9.01 15.14
Llama-3.2-3B 21.78 4.86 8.28 6.12 7.28 8.44 7.82 29.07 28.07 7.77 15.46
DeepSeek-V3 62.29 52.81 26.79 35.55 55.38 27.30 36.57 56.02 35.65 10.40 17.46
GPT-4o 60.05 50.62 27.88 35.96 54.16 28.86 37.65 59.18 30.78 9.72 15.54
Claude-3.5-Sonnet 66.43 59.06 31.90 41.42 63.06 33.25 43.54 62.32 28.59 8.90 14.41
Table 3: The experimental results of the four LLMs in Review Composition. “$S_{t}$” refers to the title retrieval rate as defined in Eq 4. While “P” represents the Precision, “R” denotes the Recall. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
  Discipline Citation Count Precision
DeepSeek Claude DeepSeek Claude
Biology 763 678 55.55 58.00
Mathematics 2288 1984 60.00 62.22
Physics 894 652 47.62 56.19
Chemistry 1334 1079 43.14 43.80
Social Science 1321 1151 46.80 56.70
Technology 904 748 44.88 49.01
Table 4: Average citation counts and reference precision across disciplines.
@@ 4.3 Analyze LLM-Generated References from Different Dimensions
In both Reference Generation and Review Composition tasks, we ask LLMs to generate references. The overall performance was discussed in the previous section. In this section, we provide a detailed comparison of the accuracy of LLM-generated references across various dimensions, as shown in Figure LABEL:f2. As seen in Figure LABEL:f2a, Claude-3.5-Sonnet demonstrates a clear advantage over other models across all dimensions in the Reference Generation task. Additionally, the accuracy of reference generation in the Reference Generation task for Claude-3.5-Sonnet, GPT-4o, and Qwen-2.5-72B follows a consistent trend across all dimensions, with the highest accuracy observed in the title dimension. Accuracy for journal name, page, and author is also relatively high. However, DeepSeek-V3 performs worse in the author dimension compared to the other dimensions. In contrast, Llama-3.2 demonstrates higher accuracy in the page and author dimensions than in other dimensions. However, overall, Llama-3.2-3B does not exhibit a competitive advantage in reference generation accuracy.
Next, we examine the accuracy of LLM-generated references across various dimensions in Review Composition, as shown in Figure LABEL:f2b, and comparing it with Figure LABEL:f2a, we observe improvements across all dimensions for Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B, with particularly obvious gains in the author dimension. The possible reason is that, in the generated text, the LLMs tend to cite the first author’s name, which may lead the models to place more emphasis on this dimension. Notably, the accuracy of DeepSeek-V3 and GPT-4o in certain dimensions approaches or even exceeds that of Claude-3.5-Sonnet. However, the performance of LLaMA-3.2-3B remains suboptimal.
@@ 4.4 Cross-Disciplinary Analysis
First, we observe that in the Reference Generation task, as shown in Figure LABEL:f4a, almost all models exhibit the highest precision in the Mathematics discipline and the lowest precision in the Chemistry discipline. To validate these differences, we conduct one-way ANOVA tests for each LLM across five disciplines. Significant differences are found for all models except Llama-3.2-3B. Detailed ANOVA results are reported in Appendix G.
@@
Secondly, as shown in Figure LABEL:f4b, the NLI scores evaluated by TRUE in the Abstract Writing task indicate that all models perform the worst in Social Science. GPT-4o performs best in Technology, while Claude 3.5-Sonnet achieves the highest performance in Biology. One-way ANOVA tests reveal significant differences across disciplines for all models. See Appendix G for detailed results.
@@
Thirdly, we examine the references precision of each model across five disciplines in the Review Composition task, as illustrated in Figure LABEL:f4c. It is evident that the precision of Claude 3.5 Sonnet, DeepSeek-V3, and GPT-4o is significantly higher than that of Qwen-2.5-72B and LLaMA-3.2-3B across all disciplines. Furthermore, Claude 3.5 Sonnet, DeepSeek-V3, and Qwen-2.5-72B exhibit the highest precision in Mathematics, while GPT-4o performs best in Social Science. ANOVA tests confirm significant differences across disciplines for all models (see Appendix G).
@@
Finally, we observe the KPR metric across different disciplines in Review Composition, as shown in Figure LABEL:f4d. The results from the figure indicate that the differences between models—Claude-3.5-Sonnet, DeepSeek-V3, GPT-4o, and Qwen-2.5-72B—are not significant across various disciplines, a finding that is also supported by statistical tests (see Appendix G).
Citation Frequency and Precision Across Disciplines. Additionally, we report statistics on the citation frequency of correctly generated references by LLMs in the Reference Generation task, as shown in Table 4, using Claude-3.5 and DeepSeek-V3 as examples. The data indicates that the references generated by the LLMs are highly cited, which might be due to their frequent presence in online sources, making them more likely to appear in the LLMs training datasets. As a result, LLMs tend to generate more accurate metadata (e.g., author, year) for these well-known references.
Furthermore, when analyzing different disciplines, we observe that the Mathematics discipline has the highest precision, and the relevant references generated for Mathematics also have the highest citation count. We compute the correlation between citation precision and average citation counts, finding that the correlation coefficient for Claude-3.5 is 0.4, and for DeepSeek-V3 it is 0.51, indicating a positive relationship between the two.
@@ 5 Conclusion
In this paper, we evaluate the ability of LLMs to write literature reviews. However, several limitations remain:
First, instead of evaluating the generated reviews from conventional perspectives such as fluency or topic coverage, we primarily compare LLM-generated results with human-written ones. As such, our current evaluation metrics may not be comprehensive. In the future, we plan to incorporate additional aspects of review quality to improve the completeness of our evaluation. These may include the coverage of cited works (i.e., whether the review offers a comprehensive overview of the relevant field) and the coherence of the overall structure (i.e., whether the review is organized in a way that facilitates information-seeking).
Second, there is a possibility that our test data overlaps with the training data of the LLMs. When we initiated this study in August 2024, the dataset from the Annual Reviews website had not yet been updated to include 2024 articles, so we relied on the complete 2023 dataset. To mitigate potential data leakage, we plan to deploy a leaderboard on Hugging Face to continuously evaluate the performance of various LLMs in literature review writing, with real-time updates to the test dataset. However, due to the rapid iteration of LLMs, data leakage cannot be completely ruled out. That said, our experimental results—particularly those related to reference generation—show that all models still perform poorly. If data contamination were present, the actual scores would likely be lower than those reported. This reinforces, rather than undermines, our conclusion that significant challenges remain in using LLMs for literature review generation.
Additionally, when processing LLM-generated outputs, we frequently encountered abbreviated author names and journal titles. Although we have taken care to address these issues thoroughly (see Appendix E), minor discrepancies may still exist.
Finally, to verify the precision of LLM-generated references, we primarily used Semantic Scholar as our auxiliary tool. Although we also experimented with Google Scholar, its lack of an accessible API led us to rely on the freely available Semantic Scholar API for consistency and ease of access. However, this may have resulted in incomplete reference retrieval.
Ethics Statement
The human evaluations conducted in this study were carried out by members of the research team. No personal or sensitive information was collected, and all participants were fully informed of the purpose of the evaluation. Therefore, the study does not raise any ethical concerns.
@@ Appendix C Comparison of LLM-Cited and Human-cited References from Different Dimensions
We provide a more detailed comparison of the LLM-cited and human-cited references across various dimensions. As shown in Figure LABEL:f3, for Reference Generation, we observe that the overlap rate is higher in the “Title” and other numerical dimensions, while the overlap rates for the “Journal” and “Author” dimensions are relatively lower. For Review Composition, Claude-3.5-Sonnet and GPT-4o exhibit a higher overlap rate on the “Author” dimension compared to Reference Generation. This trend is consistent with the findings in Figure LABEL:f3, as the citation of author names in the literature for Review Composition leads to the generation of more accurate author information.
@@ Appendix D Discussion
When comparing the references generated by the models in Reference Generation and Review Composition, we find that in Review Composition, nearly all models generate more accurate references. This suggests that LLMs cite references during the writing process, which improves the authenticity of the references. Moreover, the inclusion of the first author’s name in the generated context also enhances the accuracy of the author dimension.
@@
Appendix E Data Processing Strategy
Author name and journal title variations often pose challenges when aligning LLM-generated references with articles from Semantic Scholar. To address this, we adopt the following normalization strategies:
Author names. When comparing author names between LLM-generated references and candidate articles, if an exact match is not found (e.g., “John Smith”), we consider common variants such as “Smith, John”, “Smith, J.”, or “J. Smith” to account for different citation formats.
Journal titles. For journal names like Journal of Chemical Physics, we incorporate standard abbreviation forms (e.g., J. Chem. Phys.) based on widely used abbreviation conventions. Nonetheless, certain non-standard or ambiguous cases may still be unmatched.
Appendix F LLM Output Example
Here is one example of the output generated by Claude-3.5-Sonnet for the literature review composition task.
{"Literature Review": "CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules. This review explores the technical developments and application prospects of CRISPR-based biosensing strategies, highlighting their potential to revolutionize diagnostics and molecular detection.The core biochemical properties of CRISPR-Cas systems provide the foundation for their application in biosensing. These properties include diverse reaction temperatures, programmability in design, high reaction efficiency, and recognition specificity. Chen et al. (2018) demonstrated the versatility of CRISPR-Cas9 in detecting DNA targets across a wide temperature range, from 20°C to 60°C, enabling its use in various environmental conditions. The programmability of CRISPR systems allows for the easy design of guide RNAs (gRNAs) to target specific sequences, as shown by Gootenberg et al. (2017) in their development of the SHERLOCK (Specific High-sensitivity Enzymatic Reporter unLOCKing) platform for nucleic acid detection.The high reaction efficiency of CRISPR-Cas systems contributes to their sensitivity in biosensing applications. Li et al. (2019) reported a CRISPR-Cas12a-based assay capable of detecting attomolar concentrations of DNA targets, demonstrating the potential for ultra-sensitive detection. The recognition specificity of CRISPR systems allows for precise discrimination between closely related sequences, as evidenced by the work of Myhrvold et al. (2018), who developed a CRISPR-based diagnostic platform capable of distinguishing between Zika and Dengue virus strains with single-base resolution.Recent efforts have focused on improving these core properties to enhance the performance of CRISPR-based biosensors. For example, Harrington et al. (2018) engineered Cas12a variants with improved activity and specificity, expanding the range of targetable sequences and reducing off-target effects. Similarly, Gootenberg et al. (2018) developed an enhanced version of the SHERLOCK system, incorporating lateral flow readouts and multiplexing capabilities, further improving its utility in point-of-care diagnostics.One of the key areas of technical development in CRISPR-based biosensing is improving sensitivity and quantification capabilities. Wang et al. (2020) introduced a CRISPR-Cas12a-based digital PCR method that achieved attomolar sensitivity and precise quantification of target nucleic acids. This approach combines the specificity of CRISPR recognition with the quantitative power of digital PCR, opening new avenues for highly sensitive and accurate molecular diagnostics.Multiplex detection is another important aspect of CRISPR biosensing development. Ackerman et al. (2020) demonstrated a CRISPR-Cas13-based multiplexed assay capable of simultaneously detecting and distinguishing multiple SARS-CoV-2 variants, showcasing the potential of CRISPR technology in rapid and accurate pathogen identification. The ability to detect multiple targets in a single reaction not only saves time and resources but also provides more comprehensive diagnostic information.Efforts to develop convenient one-pot assays have led to simplified and more user-friendly CRISPR-based detection methods. Joung et al. (2020) reported a one-pot, CRISPR-Cas12-based assay for SARS-CoV-2 detection that could be performed at a single temperature, eliminating the need for complex thermal cycling equipment. This advancement brings CRISPR-based diagnostics closer to point-of-care applications, where simplicity and ease of use are crucial.The creation of advanced sensors has further expanded the capabilities of CRISPR-based biosensing. Bruch et al. (2019) developed a CRISPR-Cas13a-based electrochemical sensor for nucleic acid detection, demonstrating the potential for integrating CRISPR technology with electronic readout systems. This approach offers the possibility of developing portable, electronic CRISPR-based diagnostic devices for rapid and sensitive detection in resource-limited settings.While CRISPR-based biosensing has shown great promise, there are still obstacles to its widespread commercial application. These include challenges in sample preparation, the need for standardization across different CRISPR systems and assay formats, and regulatory considerations for diagnostic use. Addressing these issues will be crucial for the successful translation of CRISPR biosensing technology from the laboratory to real-world applications.Looking ahead, the future of CRISPR-based biosensing appears bright, with numerous opportunities for further development and application. Potential directions include the integration of CRISPR biosensors with microfluidic devices for automated sample processing and analysis, the development of CRISPR-based wearable sensors for continuous health monitoring, and the application of machine learning algorithms to improve the design and performance of CRISPR biosensing systems.In conclusion, CRISPR-based biosensing strategies represent a rapidly evolving field with immense potential to transform molecular detection and diagnostics. As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.", "References": [{"title": "CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity", "authors": "Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "436", "last page": "439"},{"title": "Nucleic acid detection with CRISPR-Cas13a/C2c2", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.", "journal": "Science", "year": "2017", "volumes": "356", "first page": "438", "last page": "442", "DOI": "10.1126/science.aam9321"}, {"title": "CRISPR-Cas12a-assisted nucleic acid detection", "authors": "Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin", "journal": "Cell Discovery", "year": "2019", "volumes": "5", "first page": "1", "last page": "4"}, {"title": "Field-deployable viral diagnostics using CRISPR-Cas13", "authors": "Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "444", "last page": "448"}, {"title": "Enhanced proofreading governs CRISPR–Cas9 targeting accuracy", "authors": "Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.", "journal": "Nature", "year": "2018", "volumes": "563", "first page": "621", "last page": "625"}, {"title": "Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng", "journal": "Science", "year": "2018", "volumes": "360", "first page": "439", "last page": "444"}, {"title": "Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay", "authors": "Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua", "journal": "Nature Communications", "year": "2020", "volumes": "11", "first page": "4711", "last page": "4711"}, {"title": "Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing", "authors": "Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P. J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem", "journal": "New England Journal of Medicine", "year": "2020", "volumes": "383", "first page": "1492", "last page": "1494"}, {"title": "CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2", "authors": "Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.", "journal": "Angewandte Chemie International Edition", "year": "2019", "volumes": "58", "first page": "17571", "last page": "17575"}, {"title": "Scalable and robust SARS-CoV-2 testing in an academic center", "authors": "Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared", "journal": "Nature Biotechnology", "year": "2020", "volumes": "38", "first page": "927", "last page": "931"}]}
Appendix G ANOVA Test Results
We report the p-values from one-way ANOVA tests across disciplines for each model and task:
  • Reference Generation task: Claude-3.5-Sonnet (p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B (p<.0001), Llama-3.2-3B (p=0.065).
  • Abstract Writing task (NLI scores): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<0.05), GPT-4o(p<.01), Qwen2.5-72B (p<.001), Llama-3.2-3B (p<.0001).
  • Review Composition (Reference accuracy): Claude-3.5-Sonnet(p<.0001), DeepSeek-V3(p<.0001), GPT-4o(p<.0001), Qwen-2.5-72B(p<.0001), Llama-3.2-3B(p<.001).
  • Review Composition (KPR metric): Claude-3.5-Sonnet (p=0.46), DeepSeek-V3 (p=0.23), GPT-4o (p=0.18), Qwen-2.5-72B (p=0.10), and Llama-3.2-3B (p<0.001).